OpenStudy (anonymous):

Use the FOIL method to find the product for the following (7x^8 - 3)(x^9 - 5)

OpenStudy (saifoo.khan):

\[7 x^{17}-3 x^9-35 x^8+15 \]

OpenStudy (anonymous):

Thank you!!! I've got 3 more questions!!!

OpenStudy (saifoo.khan):

Post them!

OpenStudy (anonymous):

\[3w-12/35\div w-4/77w\] divide

OpenStudy (saifoo.khan):

this?\[\frac{3w-12}{35} \div \frac{w-4}{77w}\]

OpenStudy (anonymous):

yes please

OpenStudy (saifoo.khan):

\[\frac{3w-12}{35} \times \frac{77w}{w-4}\]

OpenStudy (saifoo.khan):

\[\frac{33w}{5}\]

OpenStudy (saifoo.khan):

which week are u in? 3-4?

OpenStudy (anonymous):

Thank you, I've posted one of the other questions on the board

OpenStudy (anonymous):

week 8, I'm going back over several things trying to get a grasp one it before the final

OpenStudy (saifoo.khan):

Awesome. i hope u do good. Final has 66 questions. Be ready! =D

OpenStudy (anonymous):

thank you so much for your help!!!

OpenStudy (saifoo.khan):

You welcome. =)